If a charge of magnitude 6.5x10^17C is being held in a place 2x10^-8m from a charge of -8x10^-19C Which is also being held in place. What is the electric potential of the system

1
Expert's answer
2022-03-01T11:11:33-0500
"W=k\frac{q_1q_2}{r}\\
=9*10^9*\frac{6.5*10^{-17}*(-8*10^{-19})}{2*10^{-8}}=-2.34*10^{-17}\:\rm J"